OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)?

OpenStudy (anonymous):

the equation of the line, (Y-y1)=m(x-x1) where,(x1,y1) is the given point and m is the slope

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) here, (x1,y1)=(3,-1) (x2,y2)=(-1,5)

OpenStudy (anonymous):

first find out the value of m i.e. slope by puttting these values. adn then put it all in above equation.

OpenStudy (anonymous):

whats the answer